OpenStudy (anonymous):

square root of 19. its easy to everyone, but i still fail with it.

OpenStudy (anonymous):

are you supposed to just write this in simplest form?

OpenStudy (unklerhaukus):

\[√19\]

OpenStudy (unklerhaukus):

more than 4, less than 5...

OpenStudy (anonymous):

well im in the middle of getting the area of triangles and everytime i hit the square root button on the calculator it shows 4.6436346 random numbers.

OpenStudy (anonymous):

thats because the square root of 19 is not a nice, pretty whole number. it is a nonterminating, non repeating decimal. Your calculator can only round to about 10 places. When you don't get a whole number for an answer you use the square root symbol

OpenStudy (anonymous):

so i just but sqrt19?

OpenStudy (anonymous):

yeah, you just leave it like that, unless the directions tell you to round the answer to a certain place. Thats when you use the square root button on the calculator

OpenStudy (anonymous):

do u know how to get the area of triangles?

OpenStudy (anonymous):

if you know the height it is \[\frac{1}{2}bh\]

OpenStudy (anonymous):

if you know the lengths of two sides and the angle between them it is \[\frac{1}{2}ab\sin(\theta)\]